link for news highlights about the STEREO mission.
Below, we list
additional entries concerning the STEREO instruments and data.
November 17, 2015
The STEREO Ahead observatory has returned to full operations on the main lobe
of the high gain antenna. The low rate data stored on the recorder since the
start of side lobe operations has been read out, and the recorder has been
restored to its original configuration. All the instruments have resumed
normal operations. The IMPACT SWEA and STE detectors will be powered on and
brought back up to full operations over the next few days.
November 9, 2015
The STEREO Ahead observatory has returned to using the main lobe of the high
gain antenna. From this point on, space weather beacon data will be available
again from the NOAA antenna partners between DSN tracks. The mission
operations team will now start to download the low rate data that was stored on
the recorder since the start of side-lobe operations on August 20, 2014. Once
this is completed, the recorder will be reconfigured, and the instruments will
be returned to their prime science mode. In the meantime, the instruments will
continue to produce telemetry at the side-lobe rates for the next few days.
See the planning page for more details about the
July 16, 2015
The spacecraft guidance-and-control system was returned to normal operations
when reaction wheel #3 was power cycled and resumed responding to torque
July 15, 2015
The IMPACT team reports that all of their instruments are now working and
producing science data, with the exception of SWEA and STE which won't be
turned on until the spacecraft returns to the main lobe of the high gain
antenna in November. Because of the current difficulties with commanding due
to solar scintillation effects at such a small separation angle from the Sun,
the SIT instrument is using the original table set that the instrument was
launched with. This table set is known to have some calibration problems, but
this is considered to be better than having no data at all. He intensities are
about a factor of five too low, while the H, O, and Fe intensities are
approximately correct (within a factor of two). The IMPACT team plans to
upload the most current table set when the command link is more stable, which
is expected to be in several week's time.
July 14, 2015
JPEG images from the SECCHI EUVI and HI2 telescopes are now appearing on the
SSC website. Images from the other telescopes will be made available as soon
as enough data have been collected to derive instrumental backgrounds. The
SECCHI team is also working through the issue of updating the pointing
information in the images to take into account the changed attitude of the
spacecraft after solar conjunction. The headers in the COR1 and COR2 FITS
files have been corrected, and work is proceeding on the other telescopes.
Incorrect EUVI pointing information is currently interfering with the
production of heliographic maps. It is expected that this will be corrected
The IMPACT team reports that MAG, SEP, and HET are producing science. Regular
science data from SIT-A from July 13 after the HV was ramped up showed that the
instrument MCPs were triggering normally, and events were being analyzed.
However, the on-board rate computation was returning 0 counts in all boxes, due
to the dropouts in the table loads experienced on July 13. Given the current
issues with dropped commands due to solar interference, it was decided to
reboot SIT-A, since this loads the original table set from SEP central. This
table is not optimally calibrated, but approximately calibrated data is better
than no data at all. The team's plan is to upload the up-to-date table set
when the STEREO commanding link is more robust, perhaps on August 1st, when the
SPE angle is at 3.0 degrees.
The PLASTIC MCPs have been brought up to 2590V, which was the original
operating voltage at the beginning of the mission. The PAC was brought to
-18kV, monitored for stability, and then brought to the full mission operating
value of -20kV. The instrument was then put into mode 3 to collect monitor
rates. The entrance system is not yet enabled.
July 13, 2015
On Saturday, Jul 11, the SECCHI instrument was powered back on. The team
verified with the EUVI telescope that the spacecraft was correctly pointed, and
re-opened the COR1 and COR2 doors. All the telescopes are operating normally,
but some ground software issues are currently interfering with the production
of JPEG images for the SSC website. The IMPACT team reports that HET is now
producing science. Due to solar interference, efforts to patch LET have been
slowed. The SIT-A high voltage ramp-up was completed today, and the analog HK
values from the instrument are at nominal levels. The largest of the lookup
SIT tables has not been successfully reloaded, apparently due to ground GSE
problems, but the smaller SIT tables have now been loaded. PLASTIC continues
to ramp up their high voltages.
July 10, 2015
The IMPACT team reports that MAG and SEPT are producing science. Due to
communication interference, efforts to patch LET have been slowed and will
continue on Monday. HET is producing some science, but a table load needs to be
completed on Monday to get back to its nominal state. SIT has a bit more
configuration to do as well plus its high voltage needs to ramp up so it will
be sometime next week before it is producing science. All these instruments
appear to be healthy. No anomalous behavior has been seen after the long
July 9, 2015
Post-solar-conjunction instrument recovery continued with the powering on of
the IMPACT and PLASTIC instruments. The IMPACT MAG, SIT, HET, LET and SEPT
subsystems have been powered on. SWEA and STE will remain off until the
spacecraft returns to the main lobe of the high gain antenna later this year.
MAG has already started producing science data. The SEP suite
is expected to be in science mode sometime on July 10, except for SIT which
will take several more days to bring back to full operating mode.
July 8, 2015
The STEREO Ahead spacecraft was returned to three-axis stabilized mode today,
and the operations team has started the process of returning the spacecraft to
normal operations. The scientific instruments are expected to be turned on
over the next few days. The S/WAVES radio instrument has remained on all
throughout the superior solar conjunction period, and has started to return
data again. More information on the status of STEREO Ahead will be
provided as it becomes available.
June 4, 2015
A simulation of one possible model of the events that led up to the loss of
contact with STEREO Behind is available here.
March 24, 2015
The STEREO Ahead was reset into its "safe mode"
16:00 UT on 24 March 2015 in preparation for superior solar conjunction.
Normal operations are expected to resume in July 2015. Please see
this link for more
March 20, 2015
The IMPACT, PLASTIC, and SECCHI instruments on STEREO Ahead were turned off at
16:00 UT on 20 March 2015 in preparation for superior solar conjunction. They
are expected to resume operation in July 2015.
March 18, 2015
November 19, 2014
November 7, 2014
A new web page has been put together describing the recent
loss of contact with the STEREO Behind
October 7, 2014
Communications with the STEREO Behind spacecraft were interrupted on October 1,
immediately after a planned reset of the spacecraft performed as part of a test
of solar conjunction operations. The cause of the anomaly is not yet known,
though a sensor anomaly in the guidance and control system is suspected.
Attempts to recover the spacecraft are continuing.
September 3, 2014
A massive proton storm caused by a large flare on the far side of the Sun hit
both STEREO spacecraft starting September 1, 2014. The high energy particle
fluxes were so high that the star tracker reset on both spacecraft reset within
minutes of each other. The star tracker on Behind returned to full operation
after 4.7 hours, and on Ahead after 18.7 hours. Both star trackers are now
The storm was so massive that many of the SECCHI images during the storm do
not pass their quality check, and therefore do not appear on the daily browse
pages. Other images from the storm period that do make it onto the web pages
are heavily contaminated by the proton storm, and are highly distorted. This
is particularly true for the HI-1 and HI-2 images taken during the height of
August 20, 2014
A new procedure has been developed for processing HI-2 images during side lobe
operations. The old procedure depended on having pairs of images 2 hours
apart, in a process known as "running difference". This worked fine when we
were getting 24 hour coverage. However, now that the Ahead spacecraft is
operating on one of the side lobes of the high gain antenna, we only get a few
hours of real-time data per day, during the periods when the spacecraft is in
contact with one of the 70 meter ground stations. With such restricted
telemetry, only a small number of HI-2 images has a preceeding image to form
the running difference. The new technique uses nearby images to generate a
background image for subtraction, and allows all the images to be
processed. Experimentation shows that coronal mass ejections still show up in
these images, though not quite as clearly as in the running difference
technique. Behind data will continue to be processed as running differences
until that spacecraft also starts using the antenna side lobe at the end of
November. For more information about the STEREO side lobe operations, please
August 20, 2014
The first telemetry downlink from the STEREO Ahead spacecraft during the
side-lobe operations period has been carried out successfully. Beacon data was
collected from all the instruments, and processed for the real-time image and
plots pages. There are some minor issues which have been identified in the
ground processing of the SECCHI images in this new mode of operations, but
they should be corrected soon. The spacecraft performed as expected. For more
information about the STEREO side lobe operations, please see this
August 19, 2014
The STEREO Ahead spacecraft has been reconfigured for side lobe operations.
The solid state recorder has been played back, and repartitioned to store
spacecraft housekeeping and in situ beacon telemetry for eventual
playback once normal operations are resumed in January 2016. The high gain
antenna has been moved to the first side lobe position; the first pass using
this configuration will be on Wednesday, August 20. See this
important notice for more information.
August 15, 2014
The change in COR1 processing implemented on August 14 did not work out as
expected, and the software has been returned to its previous behavior.
However, the initial backgrounds for the most recent COR1-Ahead data have been
August 14, 2014
Because of the difficulty in deriving good instrumental backgrounds at this
stage of the mission, the strategy for generating the browse images for the
COR1 telescopes has changed. In the old procedure, regions where the
instrumental background was overestimated would end up completely dark. The
new procedure should correct this problem, but with the side affect that some
areas in the outer corona may now appear too bright. However, it was felt that
this was preferable to completely losing visibility in parts of the image.
It's anticipated that after sufficient time has elapsed to provide better
backgrounds, and the images have been regenerated, the difference between the
two methods will be minimal.
July 11, 2014
The test of operating the STEREO Ahead spacecraft during the upcoming period of
reduced operations has been successfully
concluded. Recovery of the science instruments after being off for the last
several days has begun.
June 26, 2014
Unexpectedly high temperatures in the high gain antenna feed horns on both
STEREO spacecraft will require corrective action in the coming months that will
severely limit science operations. The high temperatures are being caused by
the small angle between Earth and the Sun as seen from each spacecraft. In
other words, pointing the antenna at Earth is putting too much solar heat on
the antenna feed horn. To bring down the feed horn temperature, and preserve
the spacecraft for years to come, the antennas will be pointed off at an angle
from both Earth and the Sun, so that less heat will fall on the feed horns.
Communication will still be possible using one of the antenna side lobes, but
the telemetry rate will be extremely low. What instrument operations
can be supported while the antennas are off-pointed is still being studied.
A test of operating the STEREO Ahead spacecraft in this antenna off-pointed
mode will be conducted during the week of July 6-12, 2014. All science
instruments will be turned off during this time, and instrument recovery
operations are expected to take several days.
More information about the upcoming period of reduced operations can be found
on this page, which will be updated as
more information becomes available.
May 30, 2014
A new version of the SECCHI/HI Level-2 product is
available on the STEREO Science
Center website. These data have have actually been partially available for a
while now, but announcement was delayed until the ingestion of the files was
mostly complete. At the time of this writing, some data from the first year of
the mission is still being ingested, but the data should be complete from 2008
The new data are available in three different units: detector data numbers per
pixel per second (DN/Pixel/Sec), Mean Solar Brightness (MSB), and the
astronomical diffuse magnitude scale known as S10. The MSB and S10 data have
had the diffuse brightness correction applied. The
file naming convention has been updated
for these data.
January 6, 2014
The Inertial Measurement Unit (IMU) on STEREO Behind failed at 9:46 UT on
Sunday, January 5, putting the spacecraft into coarse point mode. Recovery
operations are currently being carried out. All the instruments continue to
take data. However, the COR1 and COR2 doors are closed.
Update: Fine pointing mode has been restored on STEREO Behind. The COR1
and COR2 doors have been reopened.
August 13, 2013
July 22, 2013
The STEREO Behind spacecraft has been completely recovered, and is operating
nominally. The cause of last week's transponder anomaly is still under
July 19, 2013
The STEREO Behind spacecraft is currently in a low-rate telemetry mode. The
mission operations team is diagnosing the problem. More information will be
posted as it becomes available.
Update:The The transponder on the STEREO Behind spacecraft has been
recovered back into 120 kbps mode. The remaining recovery activities,
including activating the beacon rate between passes, will be completed on the
next pass. The rest of the spacecraft has been operating normally the entire
time, and the instruments have been collecting data.
February 14, 2013
A new web page has been put together describing the possible observations of
Comet ISON by STEREO
in October-December 2013.
December 4, 2012
The mode of operations for the COR1 telescope on STEREO Behind was changed back
to the normal sequence of 0, 120, and 240 degrees.
November 29, 2012
Starting November 29, the mode of operations on the COR1 telescope on STEREO
Behind was changed to take images at polarizer angles of 60, 180, and 300
degrees instead of the previous 0, 120, and 240 degrees. This was done in an
attempt to correct occasional position errors that had developed in the 240
degree images. Changing the polarizer angles introduces subtle effects in the
instrumental background. It will take some time for sufficient data to be
collected to create new background files. In the meantime, provisional
background files have been formed based on the older data. As a result, the
COR1-B images will appear noisier than usual until new background files are
November 27, 2012
Users may have noticed that there have recently been a significantly smaller
number of SECCHI images from the Behind spacecraft, and large gaps in the
IMPACT, PLASTIC, and SWAVES data from Behind. This is due to a shortage of
telemetry coverage caused by Deep Space Network schedule conflicts with other
missions. This situation started in mid-November 2012, and is expected to
continue for the next several months. In upcoming weeks, the Ahead spacecraft
will also start to be affected.
November 20, 2012
July 11, 2012
The daily STEREO
pages now include daily synoptic maps generated from a combination of STEREO
and SDO images, in all four EUVI wavelengths. These daily maps have been
generated back through 2010, and earlier dates will be available soon.
March 26, 2012
searchable STEREO publication database is now available. This replaces the
old database system which had to be taken offline. Authors of STEREO-related
articles in the refereed literature are encouraged to submit them for inclusion
in the database once they are accepted for publication. It would be
particularly helpful to include the DOI, volume, and page number.
December 14, 2011
A new running difference browse product is now available for the SECCHI EUVI,
COR1, and COR2 telescopes. These running difference images are available on
the individual daily browse pages, and on
the latest images page, and should be
available soon via the image search tool.
Currently, running difference images are available only for dates starting
November 1, 2011, but earlier dates will be filled in over time.
July 22, 2011
The power outage earlier reported on the home page for July 22-23 has been
July 20, 2011
The software used to generate the SECCHI browse images has been changed to more
closely reflect the actual data. In the previous version of the software, when
the JPEG images being generated were larger than the original images, the data
were interpolated to the larger size. The new version of the software resizes
the images without interpolation. The result is a less-smooth, more
"pixelated" image, but which is more representative of the actual data.
Particularly affected are the beacon images. The smoother beacon images
produced by the previous version sometimes confused people into thinking that
they saw structures which were not supported by the actual data.
May 18, 2011
The SECCHI EUVI, COR1, and COR2 telescopes on STEREO Behind picked up an
interference pattern on 9 May 2011 around 17:20 UT, thought to be due to a
single event upset to the Camera Electronics Box (CEB). The CEB was restarted
to fix the problem on 12 May 2011; however, in the meantime the background
images for COR1 and COR2 had become corrupted, preventing the team from
immediately realizing that the problem had been solved. New backgrounds are
now being produced, and the browse images are being
regenerated as corrected backgrounds become available.
April 5, 2011
A new "Downlink Performance" data product is now available as part of the
Ancillary Data set.
January 13, 2011
The IMPACT/PLASTIC Data Processing Unit on Behind was autonomous switched
off on 8 January 2011 due to unexpectedly high current. The IMPACT and
PLASTIC instruments were switched back on, and all the instruments are now back
into normal operating mode.
December 17, 2010
The SECCHI team is now making available movies made up from the last seven days
of data. These movies are available from their
page, under the sidebar heading "Current Weekly Movies". Note that each STEREO
spacecraft is now within the field of view of the COR2 telescope on the other
December 16, 2010
The PLASTIC team has instituted a new data browse product called "Protons
Projected to Earth" on their
page. These overplot the STEREO A, B and Wind proton data sets, with a simple
time translation assumed for A and B. The purpose is to demonstrate how
STEREO-B data can be used as an advance proxy for L1.
December 1, 2010
The remaining problems with connectivity between the STEREO Science Center and
the Mission Operations Center after the move on November 29 have been resolved.
All SSC services should now be functioning normally.
November 30, 2010
The STEREO Science Center has been successfully moved to its new location, and
the website and beacon services are back up and running.
November 5, 2010
Level-2 data are now available.
September 8, 2010
August 11, 2010
July 8, 2010
We are now making available a prediction of the
state of the Sun during the upcoming total solar eclipse on 11 July 2010, based
on what STEREO Behind saw at the same heliographic longitude on 6 July 2010.
The problem with accessing recent AIA data for the STEREO
synoptic maps has been resolved.
July 7, 2010
We are currently having problems accessing recent AIA data for the STEREO
synoptic maps. The VSO and AIA teams are working on the problem, and we hope
to have it resolved as soon as possible.
July 2, 2010
June 24, 2010
SDO/AIA images are now being used to supply the "Earth-view" in the STEREO
Latest Images and
Daily Browse Images pages for dates after April 16,
2010. SOHO/EIT images are used for earlier dates, or if the SDO/AIA data are
unavailable. Since AIA does not observe the Fe XV line at 284 Angstroms, the
similar Fe XIV image at 211 Angstroms is used instead. Note that the SDO/AIA
team has chosen to display their images in different color tables than were
used for the SOHO and STEREO projects.
June 22, 2010
The STEREO heliographic maps on our
Images page now include SDO/AIA data for the Earthward-facing part of the
Sun. Once the two STEREO spacecraft reach 180 degree separation on 6 February
2011, the combination of images from all three spacecraft will give us a
complete view of the entire Sun.
June 17, 2010
April 20, 2010
Level-2 archive has been reorganized. The former contents have been
moved to a new subdirectory called
and new subdirectories
have been added.
January 1, 2010
The pages on image artifacts due to
beacon processing and
cosmic rays have been updated to show how these two
artefact sources interact.
November 18, 2009
A new webpage has been inaugurated showing side-by-side
comparisons between the latest
SECCHI beacon images, and the most recent GONG far-side maps. When candidate
magnetic regions appear in the GONG maps, one can look for enhanced brightness
in the corresponding area in the EUVI maps. There is still a region on the far
side of Sun which is not yet observed by STEREO, but this region will shrink
over time, and we should be able to observe the entire Sun by February 2011.
November 3, 2009
Beacon telemetry files are now
on the website for instrument teams who want to use the beacon telemetry to
fill gaps in the level 0 telemetry
November 2, 2009
The IMPACT team has removed the following directories from their Level 1 data
and the same for behind. MAG data are now to be found in the RTN and SC
directories for data in RTN and spacecraft coordinates respectively.
October 15, 2009
We have been receiving a number of questions recently about planets visible in
the STEREO telescopes. A new
STEREO Planet Finder page shows the calculated
positions of planets in the latest SECCHI beacon images. One can also use the
STEREO Orbit Tool for older data. Select either the
inner or outer solar system options to see which planets might be in the field
of view of the STEREO telescopes.
August 26, 2009
The SECCHI instrument team has instituted new software for their HI2 beacon
images. These images are now available on the
latest images and daily
browse pages. To save telemetry, only the sunward half
of the images are sent down in the low rate beacon telemetry stream.
IMPACT/PLASTIC Level-3 event lists from UCLA are now
archived within the SSC.
August 6, 2009
The IMPACT/SIT instrument has now been restored to full operation. Now all the
STEREO instruments are recovered from the system reset on August 4.
August 5, 2009
The IMPACT/SWEA high voltage on Ahead is now at its nominal value. The only
remaining detector still needing to be fully recovered is IMPACT/SIT.
August 5, 2009
PLASTIC on Ahead is back to operating mode. The high voltage supplies are
still warming up, and are expected to take several hours to reach equilibrium
temperatures. This may affect the data values during those hours.
August 4, 2009
The SECCHI telescope suite has been recovered, and is taking images again.
August 4, 2009
All the instruments have been powered back on successfully. Some beacon
telemetry is now available from IMPACT and SWAVES. Instrument recovery will
continue over the next few days.
August 4, 2009
Powering up of the instruments has begun.
August 4, 2009
The STEREO Ahead spacecraft was successfully reset today (Aug 4) to load new
guidance and control software. The spacecraft is back on the high gain
antenna, and recovery is proceeding as scheduled. Powering up of the
instruments is expected to begin soon.
August 3, 2009
The STEREO Ahead spacecraft will be reset on August 4. Complete recovery is
expected to take several days. This page, and our
status page will be updated as more information
becomes available. The expected sequence of events is as follows:
- The instrument start preparations for shutdown at 5:45 UT. High voltages
will be ramped down.
- Instruments will be powered down at 9:40 UT.
- The spacecraft will be commanded to reset around 10:30 UT.
- Powering up of the instruments is expected to commence around 15:50 UT.
- Pass ends at 17:40 UT.
Instrument recovery will continue over the next couple of days.
July 28, 2009
PLASTIC Level-2 data are now available in
July 23, 2009
The PLASTIC Level-2 directories have been reorganized. Under the directories
for each spacecraft now appear directories labeled "10min", "1hr", and "1min"
for different time resolutions. The previous data appear in the "1min"
July 15, 2009
The STEREO website has been restored. There are still a few features which are
disabled, but most things should work. In particular, the popular
image search utility is working again, as is the
orbit tool. We are endeavoring to restore the remaining
features as quickly as possible.
July 14, 2009
The STEREO/VSO interface, which went offline along with the webserver on June
22, is back online. STEREO data can now once again be retrieved either
through the web interface,
or through the IDL interface (recommended), such as through the routine
July 13, 2009
Time-elongation plots of SECCHI beacon
data are now available. These plots are designed to show the propagation of
CMEs through the telescope fields of view.
July 2, 2009
The data and browse directories have been restored to the website, but scripts
are still not functional.
June 29, 2009
The STEREO Science Center website went offline on June 22, 2009. Currently,
we're running a minimal website with some, but not all, of the functionality of
the original website. Please excuse us while we try to address the problem.
May 18, 2009
We have been informed that all instruments have now been completely recovered
from the planned spacecraft reset on May 12, 2009, including ramping up high
May 14, 2009
The STEREO Behind spacecraft was successfully rebooted on May 12, 2009 to use
the new Guidance and Control software. Recovery of the spacecraft has been
completed, and all instruments have been powered up and are operating. High
voltages on some of the instruments may still require adjustment.
May 12, 2009
Starting around 23:00 UT on May 12, 2009, the STEREO Behind spacecraft will
be reset to load a new Guidance and Control software version. Full recovery of
all the instruments is expected to take several days. Please check the
status page for updated information as it becomes
May 11, 2009
Initial processing of COR1 images was changed in an attempt to bring out more
of the solar features before the preliminary time-dependent background
subtraction has settled down to its permanent value. Later processing of the
COR1 images, at 15 and 35 days after the observation date, uses the original
This new procedure turned out to be unsuccessful, and was
May 4, 2009
The PLASTIC team now provides solar wind proton data as "condegrams", showing
five Carrington rotations at a time in a spiral format. This brings out
recurrent features (such as CIRs), and shows changes in magnitude from one
Carrington rotation to the next. These plots are available from our
instrument data page.
April 8, 2009
The SSC is now providing heliographic maps formed from the latest SECCHI Ahead
and Behind EUVI images to provide a more integrated view of the current state
of the Sun. The synoptic map, updated hourly, can be found on our
latest images page.
Note just how much of the Sun we can see now.
April 6, 2009
March 27, 2009
The IMPACT SEPT files have been reprocessed to remove the timestamp bug. The
corrected files are version V06 and above. SEPT files below version V06 should
not be used.
February 25, 2009
Updated Warning: A software bug has been discovered causing times in the
IMPACT SEPT Level 1 data to be too early by about 33 seconds. The IMPACT team
is in the process of correcting the files and appreciates your patience. They
have verified that the problem exists only for the IMPACT SEPT
level 1 data. No other IMPACT data products, including products for MAG, SWEA,
STE, LET, SIT and Beacon, are affected.
February 24, 2009
Warning: A software bug has been discovered causing times in the IMPACT
CDF files to be too early by about 33 seconds. The IMPACT team is in the
process of correcting the files and appreciates your patience. They have
verified the problem exists for the IMPACT SEPT and Beacon data files and may
also be present for several other IMPACT Level 1 data products. The times are
known to be correct for the LET Level 1 data product.
September 11, 2008
Although some technical problems may still need to be worked out, data
ingestion from NRL has resumed. The most recent data has been ingested, but
some backlogged data is still being copied over.
September 5, 2008
Because of server problems at NRL, the processing and ingestion of SECCHI FITS
data is temporarily suspended.
August 22, 2008
A new IMPACT/PLASTIC Level 1 data product including data from the 200h
housekeeping packets is now available on the SSC server for both
August 15, 2008
Ingest of SECCHI data into the Stereo Science Center was resumed on Tuesday,
August 12. The backlog of data has now all been ingested and processed into
August 8, 2008
Ingest of SECCHI data into the Stereo Science Center has been temporarily
suspended until additional disk space can be installed. Data ingestion is
expected to resume on Tuesday, August 12th, 2008. All other data products
continue to be ingested.
July 31, 2008
The first announcement
for the combined STEREO-3/SOHO-22 Workshop, "Three eyes on the Sun -
Multi-spacecraft studies of the corona and impacts on the heliosphere", is
July 22, 2008
June 5, 2008
April 14, 2008
April 10, 2008
The STEREO proposal to the 2008 Senior Review is now available from our
April 8, 2008
The PLASTIC level-1 data archived at the SSC have been moved down into a new
level1 subdirectory to make way
for new level-2 data which are now
March 20, 2008
Daily browse movies are now available from the
Data Center. This website allows one to examine side-by-side images from
STEREO Ahead and Behind, or from two SECCHI telescopes on the same spacecraft.
December 27, 2007
Normal operations have resumed on Behind, one day earlier than originally
expected. Both spacecraft are now operating as expected.
December 26, 2007
The Behind star tracker has continued to perform nominally since December
23. Should no further problems develop, Mission Operations plans to resume
normal operations on Friday's track (Dec 28).
December 23, 2007
The Mission Operations team was able today to command the star tracker on
Behind back into operational mode. The star tracker continued to operate
throughout the track. As a precaution, use of the star tracker was suspended
at the end of the track, and the spacecraft was switched from the high gain
to low gain antenna. Mission Operations will continue to staff tracks, and
proceed with caution regarding ST usage. Contingency planning will continue
as will discussion with the star tracker supplier.
December 22, 2007
Problems persist with the star tracker on Behind. The high gain antenna
continues to point to Earth, and the Mission Operations team was able to use
it to download science telemetry. All other spacecraft systems continue to
December 21, 2007
The star tracker on STEREO Behind had an anomaly at 10:10 GMT this morning.
This resulted in the spacecraft autonomously switching to the low gain
antenna. All other spacecraft systems are healthy. Diagnostic telemetry is
being collected, and corrective action is under consideration.
Science data is still being collected, but at the current low telemetry rate
it is not possible to send it to the ground. Some science data may be lost
when the onboard recorders reach their maximum capacity.
December 19, 2007
Running difference images and movies from the SECCHI Heliospheric Imager
HI-2 telescope are now available on the SSC browse and
image search pages. So far, only a few days
are available, mostly in December 2007, but more days will be populated in
the near future.
Unlike the other SECCHI telescopes, the HI-2 images are only processed from
the full resolution level-0 science data. Thus, there is a delay of several
days before any HI-2 images are available.
December 6, 2007
Data from the IMPACT STE detectors are now available on the
November 29, 2007
August 10, 2007
August 10, 2007
Images from the SECCHI HI1 telescope are now available on the
Daily Browse and Image Search
Tool pages. Because the HI1 images are derived from the level-0 science
data, they take a few days to appear on the website.
August 7, 2007
Level-1 data files are now available.
August 6, 2007
July 10, 2007
June 27, 2007
Several improvements have been made to the in-situ beacon pages at
The IMPACT magnetometer data is now in the standard Radial-Tangential-Normal
(RTN) coordinates, and total magnetic field strength has been added to the
plots. New pages have been added for solar energetic particle data.
March 22, 2007
The daily data browse pages are now available at
So far, only a few days worth of data are available, but earlier data will be
steadily added. We're also working on adding more in-situ data, as well as the
daily summary plots from SWAVES.
March 21, 2007
February 14, 2007
February 14, 2007
January 23, 2007
The Behind spacecraft flew by the Moon for the second
and last time on January 21. Both spacecraft are now in heading away from
Earth into their orbits around the Sun. The official science mission started
The baffle cover on the SECCHI Heliospheric Imager on Behind was opened on
January 11. The scientists were thrilled to find that Comet McNaught was in
the HI-1B field of view. The IMPACT SIT and SEPT covers were opened on Behind
on January 16. All instrument covers are now open.
December 22, 2006:
The A spacecraft performed its lunar flyby last
Friday and is now in heliocentric orbit. The drift rate of the spacecraft is
+21.66 degrees/year (relative to the earth), this is well within its
specification of 22+/- 2 deg/year. The B spacecraft also performed a lunar
flyby, but its trajectory has another flyby scheduled for 21 January 2007. The
B spacecraft successfully completed its final trajectory adjustment maneuver on
December 14, 2006:
All the subsystems on both spacecraft continue to
perform very well. On Monday, Dec. 11, the Ops team performed a peak power
test on the A-spacecraft. The arrays are producing slightly more power than
expected (720W vs 690W). The G&C team has started to tweak the control
algorithms with the result that mechanical oscillation of the SWAVES antennae
has completely disappeared on both spacecraft. They will continue to fine tune
the system and are planning a test which would run all the wheels through all
speeds to investigate the effect of wheel speed on jitter.
The SECCHI SCIP-B telescope doors (EUVI, COR1 and COR2) were successfully
opened Dec. 13 at 6:30am. First light images were taken and reviewed by a bunch
of very happy scientists on the east and west coasts. Performance was as
expected and comparable to the SCIP-A telescope images taken last week. In
addition, there was a successful opening of the first HI telescope door on
S/C-A. The HI images appear to be exceptional. This means that of the 10
SECCHI telescopes, 8 have successfully opened their doors and taken good
images. The HI-B door, which covers the last 2 telescopes, will open at the end
of December or early January, depending on when we verify that there will be no
more B Observatory maneuvers.
The IMPACT SEPT-NS, SEPT-E and SIT doors were deployed Dec. 14 on Ahead. In
anticipation of this exciting event, the Sun decided to become active making it
easy to confirm the door openings in the science telemetry. There were 5 door
actuations - one for SIT and two for each set of doors on SEPT-E and SEPT-NS.
The door deployments for SIT, SEPT-NS and SEPT-E on Spacecraft B will be in mid
December 4, 2006:
The SECCHI team opened the doors to the SECCHI SCIP-A
instruments and took a few first-light images. Everything went absolutely
smoothly. The three doors opened without incident. The first images look
great even though they were sent down highly compressed to keep the downlink
The EUVI was opened first and an image from each of the four quadrants looked
fine - no major tears or pinholes in the front filters. The sun is close to
being in the center of the CCD, and the resolution is beautiful.
Then we opened COR2 and took an image--also beautiful, although there is a
slight offpoint (which we expected). The exposure time is just about where we
expected, about 3-4 seconds. We then took a pB sequence of 3 images with the
result that streamers could be seen all the way to the edge of the field.
Then COR1 was opened and it was also as expected. A pB sequence shows the
inner corona nicely out to about 2.5 R, which is what we expected.
November 29, 2006:
Both the Ahead and Behind spacecraft performance
continue to perform very well with all subsystems nominal. The instruments on
both observatories continue to perform well. The SECCHI instrument team is
planning to open the doors on the STEREO-A coronagraphs next week and is
completing internal tests.
The definitive orbit solutions from FDF indicate that the STEREO A Observatory
will achieve its required drift rate without further maneuvers. The STEREO B
Observatory will require a 0.2m/s maneuver currently scheduled approximately
December 8th. Both spacecraft will swingby the moon on December 15th, with a
second lunar swingby for STEREO B on January 21st.
The guide telescope has been enabled on both spacecraft. G&C is assessing the
performance of the attitude control algorithms and will upload new parameters
November 22, 2006:
The SWAVES team is now producing daily summary plots through their website at
Follow the links to "data" and then "data plots".
Both the Ahead and Behind spacecraft performance continue to perform very well.
So far, all the planned maneuvers have been completed successfully. The
Mission Operations team is studying the current orbit of the Ahead spacecraft
to see if any more orbital maneuvers are necessary. At least one more delta-V
maneuver is planned for the Behind observatory.
The instruments on both observatories continue to perform well. They were
powered down prior to the P2 maneuvers and then powered back on successfully.
The SEP Central, HET, LET, SIT, SEPT-E, SEPT-NS detectors on IMPACT have been
successfully been powered on and checked out on both observatories. The
Caltech team was able to correlate the first 100 minutes of data from the
STEREO LET instrument on IMPACT with ACE data and the spectra were found to be
A spacecraft roll to calibrate the IMPACT Magnetometer was performed on each
observatory. Initial results show that the test was highly successful.
PLASTIC continues to ramp up their high voltage levels on their MCPs. At the
current HV levels they are starting to see counts from penetrating particles or
photons on their position detectors. The SSDs are also seeing penetrating
particles and photons.
The SECCHI team has completed their decontamination heating of the CCDs on the
COR1 and COR2 telescopes and are preparing to do the same with the EUVI and HI
November 14, 2006:
All the spacecraft subsystems are performing very
well. The Guide Telescope was enabled on the A spacecraft Sunday and
acquired the sun. The GT on the B spacecraft was enabled Monday. In both cases
the GT data looks good and the spacecraft algorithms are correctly using the
data. The spacecraft pointing improved markedly with the GT enabled. There
are more planned tests of the GT and spacecraft control algorithms in December
and more tweaking of the control algorithms.
A complete functional test of the both SECCHI suites was completed at the end
of last week and both suites are working well.
The SEP instruments on the B spacecraft were turned on for the first time
this afternoon. Initial indications are very good.
November 6, 2006:
Both spacecraft went through a 25 minute eclipse as
they went through perigee this morning. The spacecraft behaved nominally
throughout eclipse and are now on their way out to moon radius. The Ahead
spacecraft was intentionally sent to Earth Acquisition mode in order to test
the EA system. The operation of the spacecraft was completely nominal in EA
and recovery from EA to Standby was nominal. The instruments on the A
spacecraft were turned off as part of the EA test. The instruments were turned
back on and are continuing with commissioning activities.
The PLASTIC SSD high voltage has been turned on both spacecraft. The MCP high
voltages have been turn on at low voltage on both spacecraft. The MCP voltages
will be ramped up to operational voltages over the next month.
November 3, 2006 UPDATE:
This is STEREO's second week in Earth
orbit(s). It is still in it's first of four phasing orbits--prior to the first
lunar swingby. So far the planned commissioning activities have occurred
without any serious anomalies. The major deployments of solar
arrays, spacecraft separation, and high gain antennas occurred almost
immediately after launch. Since then, the SWAVES antennas on each observatory
have been deployed -- and SWAVES started producing significant science data.
This week, the significant commission events were the execution of the apogee
propulsive maneuvers on both spacecraft that raised their perigees to altitudes
that are comfortably above the atmosphere. These 'burns' were accurate to
within 0.5% -- that is considered very good considering that the observatories
have not yet been fully characterized. The other significant event was the
deployment of the four meter rigid IMPACT boom on each observatory. The MOPS
and Engineering teams performed other important activities: 1) the High Gain
Antennae were successfully exercised over their full range of motion on both
spacecraft. 2) The software team completed its assessment of observed SEU bit
flips on the C&DH and G&C computers and uplinked a patch (after a change
control board review) to correct an issue in the memory scrub software.
Currently, both observatories are operating nominally. All subsystems and
instruments on both observatories have a GREEN health status. Some IMPACT
detectors are currently saturated, possibly due to the proximity to Earth.
This effect is still being investigated. MAG is seeing higher than expected
offsets, which is correctable in software. SWAVES-Behind observes two
interference signals (100 kHz and 16 kHz) associated with the boom deployment.
With the booms deployed the spacecraft pointing has become very stable with the
pointing much less than 7 arc-seconds. We have ended our 24/7 coverage from
DSN and have gone to our nominal 3 hour tracks per day per spacecraft.
November 3, 2006:
Engineering burns were performed on the C thruster set
of both spacecraft. Both burns were completely nominal so all the thrusters
have been checked out and are working. All the spacecraft subsystems are
November 1, 2006:
Previous item should have reported that PLASTIC was
also partially powered up successfully on both observatories.
The A1 maneuver was successfully completed on both spacecraft.
The two IMPACT booms have been deployed.
The High Gain Antennae were exercised through their full ranges of motion.
October 30, 2006:
Spacecraft subsystems are performing normally. The
SECCHI instrument suites were turned on on both observatories and all telescope
subsystems passed their functional tests. The SWAVES instruments were turned
on, and both sets of antennae were successfully deployed. Parts of the IMPACT
instruments have been powered on.
October 25, 2006:
The twin STEREO observatories were launched at 8:52 p.m. EDT.